ppp:

Sylvie finds the solution to the system of equations by graphing. y = A system of equations. y equals StartFraction 2 over 3 EndFraction x plus 1. y equals negative StartFraction 2 over 3 EndFraction x minus 1x + 1 and y = x – 1

Mercury:

next time please post the full question: Sylvie finds the solution by graphing y=2/3x+1 and y=-2/3x-1 which graph shows the solution to sylvies system of equations?

Mercury:

easy way to solve this would be to simply graph both equations and see which answer choice matches your graph https://www.mathway.com/Graph